Problem Solving: How to solve:
Supppose you have 10 coins that total $0.85. Some coins are dimes and others are nickels. How many of each do you have???? Using elimination

Answer by checkley71(8403)

N--D--$
1--9--.95
2--8--.90
3--7--.85 ANSWER 3 NICKLES & 7 DIMES=$.85
-----------------------------------------------
n+d=18 or n=10-d
.05n+.10d=.85
.05(10-d)+.10d=.85
.5-.05d+.10d=.85
.05d=.85-5
.05d=.35
d=.35/.05
d=7 answer for the number of the dimes.
n=10-7=3 number of nickles.